Write an equation of a line in slope intercept form with slope 2/3 going through the point (-3,1)
1
Expert's answer
2012-09-21T08:22:15-0400
General form of the equation in slope-intercept form:
y=mx+b
Here we have m=2/3, so
y=2/3x+b
Line is going through the point (-3,1) so
1=2/3*(-3)+b
1=-2+b
b=3
Therefore we obtain
y=2/3x+3